### v1

# Effect of Overheat and Direct Flow Loading on the MEMS Bistable Flow Sensor

Abstract
We present the findings from an experimental study of a MEMS flow sensor, in which an initially curved, double-clamped bistable microbeam serves as the primary sensing element. Our research focuses on exploring how the overheat ratio, direct flow loading, and turbulence-induced vibration, affect the sequential snap-through (ST) buckling and snap-back (SB) release of an electrostatically actuated beam heated by an electric current. The sensor is fabricated from highly doped single-crystal silicon using a silicon-on-insulator (SOI) wafer. Positioned at the chip’s edge, the microbeam is exposed to airflow, enabling concurrent dynamic response measurements, with a laser Doppler vibrometer and a video camera.
Our research demonstrates that for this sensing principle, the overheat ratio can be significantly lower compared to conventional thermal sensing elements, pointing to the potential for substantial energy savings. We also emphasize the significant impact of flow angles and vibrations on the critical ST and SB voltages, which are vital for the flow sensor’s output. Additionally, we introduce the previously inaccessible experimental observation of the beam profile’s time history during the snap-through transition.
The potential impact of this research lies in the development of more robust MEMS flow sensors with enhanced sensitivity and a better understanding of their response to environmental factors, which could have broader applications in fields such as aerospace, environmental monitoring and industrial process control.
keywords:
MEMS, bistable microbeam, bifurcation-based sensor, flow sensor
[…]
[inst1]organization=School of Mechanical Engineering,addressline=Tel Aviv University, city=Tel Aviv Jaffo, postcode=6997801, country=Israel
1 Introduction
Microscale flow sensors play a key role in advancing scientific and technological endeavors. Among the various types of flow sensors, those based on microelectromechanical systems (MEMS) have become increasingly important for monitoring gas flow in industrial environments, sensing environmental flows, detecting marine conditions and analyzing biomedical flows ho1998micro . These sensors have a wide range of applications in technical fields, such as “smart skin”, the Internet of Things (IoT), and flight control (e.g., dong2017mems ; shi2020progress ).
The utilization of MEMS-based flow sensors presents several advantages, including compactness, power efficiency, high sensitivity, and compatibility with mass production requirements. Consequently, a multitude of MEMS flow sensing methods have emerged over the past three decades wang2009mems ; kuo2012micromachined ; ejeian2019design . These sensors operate based on either thermal or mechanical principles, with a majority of the previously developed flow sensors relying on the thermal principle ejeian2019design . Thermal flow sensors offer robust and straightforward manufacturing processes, linear sensor characteristics, and lack of moving parts, distinguishing them from mechanical flow sensing methods. However, their major drawbacks include high power consumption and undesired heat dissipation. Recent advancements in microelectronic device design and fabrication have aimed to mitigate these limitations of thermal flow sensors le2021fabrication . Nonetheless, mechanical flow sensors continue to play a vital role in water flow, especially for laminar flow. For example, sensors based on micro/nano-cantilevers have found diverse applications in chemical and biological sensing mathew2018review . Mechanical flow sensors often rely on drag or lift force sensing kalvesten1996integrated ; svedin1998lift ; svedin2003new , frequency monitoring bouwstra1990resonating , or artificial hair sensor designs dagamseh2011bio . Although various MEMS flow sensors are under research, specific sensing capabilities are necessary for each technical application, contingent on environmental and flow conditions.
Previous studies krylov2018IEEE ; kessler2018JMEMS introduced an alternative flow sensing principle,combining thermal and mechanical responses in a bifurcation-type sensor. These studies showed that the electrostatically and electrothermally (Joule heated) actuated initially curved double-clamped silicon microbeam exhibits bistable behavior, in other words, remaining in one of the two equilibrium position states under the same load. As a result of this bistability, the curved beam shows high responsiveness to external stimuli near two equilibrium states and has great potential for sensor implementation. This mechanical structure is pushed to its stability limits, and is highly sensitive to even the slightest changes in heat transfer caused by flow. As a result, reducing the overheat ratio (the temperature difference between the beam and the surrounding airflow) can reduce power consumption. In addition, a curved microbeam device at the chip’s edge can be used as a spatial velocity gradient sensor kessler2020flow .
The feasibility of employing such devices for airflow sensing has been confirmed through a combination of theoretical analysis and experimental testing. However, to effectively utilize the laboratory prototype in turbulent airflow measurements within practical scenarios, it is imperative to enhance our understanding of the device’s response to various influencing factors. These factors include overheat ratios (the absolute temperature to which the microbeam is heated as a multiple of the ambient temperature), the sensitivity of the microbeam to direct flow loading, as well as noise and vibrations that may arise during turbulent flow measurements, such as those over aircraft wings or wind turbine blades.
This study aims to investigate into how the sensor interacts with these factors. To achieve this, we have expanded the diagnostic system introduced in kessler2020flow to incorporate a multipoint Doppler laser vibrometer. This addition allows us to visualize the beam shape during transitions between two equilibrium states. The outcome of this extension has revealed a previously only theoretically predicted phenomenon – asymmetric snap-through motion, as described in medina2012symmetry . Furthermore, this finding sheds light on the limitations of the model presented in the paper, emphasizing the need for further refinements.
2 Materials and methods
2.1 Sensor and chip design
The bistable flow sensor based on an initially curved microbeam was introduced in  kessler2020flow ; kessler2021sampling . We briefly review the device design and its operational principle for clarity and completeness. We use an initially curved bell-shaped electrostatically actuated double-clamped beam, as shown in Figure 1, as the sensing element. The beam is fabricated from the device layer of a silicon-on-insulator (SOI) wafer, using the SOIMUMPsTM process soimumps . The beam, anchors and electrode are all made of highly doped single-crystal silicon with a resistivity of 0.02 $\Omega\cdot$cm. The beam is designed to perform an in-plane, parallel to the substrate, motion and to deflect in the $z$-direction. The nominal, as designed dimensions of the beam are given in Table 1 and definitions are according to Figure 1.
Figure 1: (a) A scanning electron microscope micrograph of the device (the scale bar is 100 $\mu m$) and device 3D schematics. The double-clamped curved beam is shown in the initial upward curved configuration. The values $V_{\mathrm{h}}$ and $V_{\mathrm{ES}}$ are the heating and electrostatic voltages, respectively (G is a ground). (b) Schematic front view of the beam, its main geometric parameters, and the airflow direction (not to scale).
Table 1: The nominal, as designed, dimensions of the beam used in the experiments.
Parameter Size ($\mu$m)
Length, $L$ $500$
Width, $b$ $25$
Thickness, $d$ $2$
Initial elevation, $h$ $2.5$
Beam-electrode distance, $g_{0}$ $10$
The operational principle of the device relies on coupled thermal and mechanical sensing. It begins with an initially curved bistable microbeam, subjected to controlled heating, using a weak electric current regulated by a constant heating voltage, denoted as $V_{\mathrm{h}}$. Simultaneously, cooling is achieved through the presence of an airflow, as illustrated in Figure 1. The voltage difference $V_{\mathrm{h}}$ is applied across the two beam anchors, with one of them serving as the ground (G in Fig. 1a).
The time-periodic electrostatic force acting on the beam is controlled by the actuating voltage $V_{\mathrm{ES}}$, applied to an electrode, as marked in Figure 1. Increasing $V_{\mathrm{ES}}$ beyond a critical snap-through (ST) threshold, denoted as $V_{\mathrm{ST}}$, triggers a dynamic shift from the microbeam’s first to the second stable equilibrium position. Conversely, reducing the voltage below the lower critical threshold, $V_{\mathrm{SB}}$, causes the beam to snap back (SB). The interaction between Joule heating and cooling via the airflow influences the curvature of the beam and the critical voltage values. Measurements of $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ are used to determine the air velocity. This device offers enhanced performance owing to its enhanced sensitivity near the ST and SB points krylov2018IEEE ; kessler2018JMEMS . Additionally, it possesses a unique capability to measure flow at two closely spaced points at two stable positions of the microbeam, allowing for the assessment of flow velocity gradients, as discussed in kessler2020flow .
2.2 Setup
The experimental setup is illustrated in Figure 2(a). During the experiments, an AFG3022C waveform generator from Tektronix is employed. It provides a voltage through channel 1, which is subsequently amplified through a signal amplifier (Trek PZD350A). This amplified voltage is then used to apply $V_{\mathrm{ES}}$  to the electrode. Additionally, to generate the electric current required for Joule heating within the beam, a constant heating voltage $V_{\mathrm{h}}$ is applied across the beam’s anchors, with this voltage being supplied by the same waveform generator through channel 2.
Figure 2: (a) Overview of the experimental setup. (b) Optical image of the experimental setup and (c) schematics of the chip from the perspective of the side camera, with the laser from the LDV, and the airflow at an angle $\alpha^{\circ}$ relative to the beam.
The real-time displacements of the beam’s center are recorded using a single-point laser Doppler vibrometer (LDV, Polytec, OFV534). The LDV focuses its laser beam on the side face of the beam, which has a width of $b=25\;\mu$m. The values of $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ are extracted from the time series recorded with an oscilloscope (DSOX2004A, Keysight), synchronously with the LDV signal.
To regulate the flow rate, we employ a pressure gauge (P-30PSIG-D/5P, Alicat) to monitor the pressure in the air supply system. A rectangular nozzle (1” High Power, EXAIR) generates a wide, uniform jet airflow directed towards the microbeam. The air supply is calibrated using a commercial anemometer (CTV 110, Kimo). Our experiments encompass a range of air velocities spanning from 0 to approximately 15 m/s. We estimate the systematic uncertainty of these measurements to be 2.8 m/s. Additionally, it is worth noting that the airflow can be applied at various angles, denoted as $\alpha^{\circ}$ in Figure 2.
The beam is strategically positioned at the chip’s edge, as depicted in Figure 2(b,c). This positioning offers the advantage of simultaneously capturing real-time beam velocity from a top view through LDV, while also enabling direct imaging of ST and SB events using a custom-built optical arrangement. The optical setup comprises a video camera (UI-6250SE-M-GL, IDS) and a horizontally mounted tube microscope (Navitar). In most of our experiments, the side camera’s use is primarily limited to the initial setup preparations, as the bulk of quantitative measurements are conducted through the LDV.
2.3 Measurement methodology
Figure 3(a) illustrates the time history of a cyclic voltage $V_{\mathrm{ES}}$ applied to the electrode (labeled in blue, left vertical axis) and the corresponding $z$ value of the beam midpoint ($x=L/2$), which is obtained using LDV (labeled in red, right vertical axis). The electrostatic voltage signal $V_{\mathrm{ES}}$ is a linear ramp with a constant rate of about 170 V per period. The analog output from the LDV controller and the input voltage signal are looped back into a digital storage oscilloscope, allowing for the synchronized recording of their time histories. The determination of $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ values occurs during the post-processing phase. This is achieved by identifying the displacement threshold transition events within the recorded data, as illustrated in Figure 3(b). These threshold events correspond to a 5 $\mu$m jump in displacement, serving as the triggers for ST and SB events.
Figure 3: (a) Cyclic time history plot of the beam displacement (red) versus the voltage $V_{\mathrm{ES}}$ (blue) between the beam and the electrode ($V_{\mathrm{h}}$=0 V). (b) Displacement as a function of the actuating voltage on the beam. (c) Electrostatic voltage (blue) and beam displacement (the other colors) as a function of time. In contrast to (a) and (b), a track filter is activated here to filter out fluctuations. (d) A close-up of the beam displacement from six different heating levels $V_{\mathrm{h}}$=0, 0.33, 0.66, 1, 1.33, and 1.66 V.
The actuation process is repeated for four cycles, operating the beam at a frequency of 400 Hz. Averaging is employed to mitigate data variability, facilitating the calculation of the mean value and its associated standard error. It is important to note that the mechanical response can be considered quasi-static due to the significantly lower frequency of the electrical signal chosen, which is of the order of $\mathcal{O}(10^{5})$ Hz compared to the fundamental frequency of the beam.
The time history of the beam center displacement is presented in Figure 3(c) for different $V_{\mathrm{h}}$ voltage levels. Crossing of dashed lines with the electrostatic voltage linear ramp signal help to indicate the values of $V_{\mathrm{ST}}$  and $V_{\mathrm{SB}}$ ,at the moments of transitions, for the case of the heating voltage $V_{\mathrm{h}}$ equal to 0 V. A closer examination of the zoom in view in Figure 3(d) reveals a noticeable shift in the time when the threshold is crossed for increasing heating voltage $V_{\mathrm{h}}$ in the range 0 $\div$ 1.66 V. This shift results in distinctly different critical electrostatic voltages $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$. We will address this shift quantitatively in the following, but we first address the bias in these measurements, due to the LDV heating of the beam.
2.4 LDV-Induced Beam Heating Bias
The specific setup explained above enables us to obtain unique results, including direct measurement of the beam’s response time history during the transition using LDV. However, the LDV system comes with a drawback known as the “loading effect,” which inadvertently heats the beam during measurements, as discussed in kassie2020effect . To assess the bias on the measured $V_{\mathrm{ST}}$  and $V_{\mathrm{SB}}$  values, we employ digital camera imaging of the beam and compare the results obtained with and without LDV. It is important to note that the image processing method for detecting ST and SB events is less precise compared to the LDV approach. Therefore, in these experiments, we extended the period to 60 seconds to achieve more accurate results. This extension was necessary because of the video camera’s lower temporal resolution, which demanded a longer measurement duration for precise observations. Consequently, the bias due to LDV heating could be larger than in the case of high-frequency measurements in our experiments, due to a significantly longer time period of the measurement.
The figures in Figure 4(a, b) show the results with “LDV on” and
Figure 4: (a) Critical snap-through voltage ($V_{\mathrm{ST}}$) and (b) critical snap-back voltage ($V_{\mathrm{SB}}$) as a function of the flow velocity $U$. Circle markers indicate measurements with LDV, while square markers indicate measurements without LDV. (c) Sensitivity of the snap-back event, a dimensionless measure of the change in voltage with respect to the reference voltage $V_{\mathrm{SB}}$ at $U=0$ m/s. Dashed lines represent a quadratic fit.
“LDV off,” represented by circles and squares. Using the LDV system increases both $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ by around 4 V and 6 V, respectively. To further quantify the influence of LDV heating, we examine the relative change in SB voltage, denoted as $\Delta V_{\mathrm{SB}}=100\times(V_{\mathrm{SB}}-V_{\mathrm{SB}}|{U=0})/V_{\mathrm{SB}}|{U=0}$ in Figure 4(c). Notably, $\Delta V_{\mathrm{SB}}$ increases from 6% to 10% at an airflow velocity of 12 m/s when the LDV is active. We will discuss this bias further in Section 3.1, where we explore how Joule heating affects the behavior of the flow velocity sensor.
2.5 Analytical model
2.5.1 Beam equilibrium
To understand the sensor’s behavior and interpret our experimental findings, we explore a reduced-order model of the double-clamped heated microbeam with initial curvature, as shown in Figure 1. While previous models in medina2014experimental ; krylov2018IEEE addressed specific aspects of sensor phenomena, such as aerodynamic loading, overheat ratio, air cooling, or dynamic effects, each model focused on distinct aspects. In this work, we present a comprehensive model that describes the combined influence of all these effects and their interactions.
The motion of the beam subjected to an axial force $N$ (positive in compression) and a distributed actuating force $f_{a}$ can be described by the following equation
[…]
Here, $z(x,t)$ and $z_{0}(x)$ are the deformed and the initial (as-designed) elevations of the beam above its anchored ends, respectively (see Figure 1(b)); $EI_{yy}$ and $EA$ are the bending and axial stiffness of the beam, where $I_{yy}={bd^{3}}/{12}$ is the second moment of area of the cross-section, $A=bd$ is the cross-sectional area of the beam, $c$ is the viscous damping coefficient, $E$ is the Young’s modulus, $\rho$ is a mass density and $\dot{(\;)}$, $(\;)^{\prime}$ denote the derivative with respect to the time and the axial coordinate $x$, respectively.
The axial force
[…]
is generated by both the beam heating of the beam gerson2010electrothermal and by a residual stress $\sigma_{res}$, estimated to be a few MPa krylov2018IEEE ; kessler2018JMEMS . In Eq. (2), $\alpha$ is the thermal expansion coefficient of Si, and $\overline{\theta}$ is the length-averaged difference between the beam temperature $T({x})$ and the ambient temperature $T_{\infty}$. In our case, we use Joule heating by applying a voltage $V_{\mathrm{h}}$ between the ends of the beam to induce axial loading.
The beam is actuated by an electrostatic force, which is parameterized by the voltage $V_{\mathrm{ES}}$, and can be described by the following equation
$f_{a}(z)=\frac{\varepsilon_{0}bV_{\text{ES}}^{2}}{2(g_{0}-z)^{2}}\left(1+0.65\,\frac{\langle g_{0}-z\rangle}{b}\right),$ (3)
[…]
We construct the single-degree-of-freedom reduced-order model associated with Eq. (1) using the Galerkin decomposition based on the following approximations of $z(x,t)$ and $z_{0}(x)$:
$z(x,t)=q(t)\phi(x),\quad z_{0}(x)=-{h}\phi({x}).\;$ (4)
where $\phi({x})=1/2[1-\cos(2\pi x/L)]$ is the base function, $q$ and $h$ are the deformed and the initial midpoint elevations of the beam, respectively. Note that $h$ is positive and the minus sign is due to the direction of the positive $z$-axis defined in Figure 1b. Substituting the representation Eq. (4) into Eq. (1), multiplying the resulting expression by $\phi(x)$, integrating by parts, and using the orthogonality of base functions in Galerkin’s formalism, we obtain the following equation expressed with non-dimensional variables
$\ddot{w}_{m}+\frac{\dot{w}_{m}}{Q}+w_{m}\left(1-\!{n}+\frac{\eta^{2}}{8}\right)-\frac{3w_{m}^{2}\eta}{16}+\frac{w_{m}^{3}}{16}=\!-{n\eta}+\frac{\beta}{\sqrt{(\gamma+\eta-w_{m})^{3}}}.\;\;$ (5)
Non-dimensional quantities used in the reduced order model are summarized in Table 2.
Table 2: Quantities used in the development of the reduced-order model.
$r=\sqrt{I_{yy}/A}$ Gyration radius
$\gamma=g_{0}/r$ Normalized beam-electrode distance
$\eta=h/r$ Initial midpoint elevation
$w_{m}=\eta+{q}/r$ Midpoint deflection
$N_{e}=4\pi^{2}EI_{yy}/L^{2}$ Euler’s buckling force
$n=\frac{N}{N_{e}}=(\sigma_{\text{res}}A+\alpha EA\overline{\theta})/N_{e}$ Normalized axial force
$\beta=\frac{\varepsilon_{0}bL^{4}V_{\text{ES}}^{2}}{8\pi^{4}r^{3}\sqrt{\gamma}EI_{yy}}(1+0.65\,\frac{\langle g_{0}-z\rangle}{b})$ Voltage parameter
$Q=4\pi^{2}/(c{L}^{2})\sqrt{EI_{yy}\rho A/3}$ Quality factor
For the case of quasi-static actuation, we can omit the terms containing time derivatives, then Eq. (5) could be solved in respect of $w_{m}(V_{\mathrm{ES}})$.
Equation (5) helps to highlight the flow sensor operational scenario. When the device is exposed to an airflow, the temperature of the beam changes, changing the average temperature difference between the beam and the air, $\overline{\theta}$. This leads to a variation in the axial force ($n$ in Table 2) and alters the critical ST and SB voltages expressed through $\beta$. To predict the $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ values and the flow, it is necessary to quantify the convective heat transfer and relate $\overline{\theta}$ to flow velocity over the beam.
For simplicity of analytical prediction, we limit our analysis to steady-state heat transfer, and neglect radiation. In this simplified case, the heat transfer equation in the beam is governed by the equation
$-(\kappa{\theta}^{\prime})^{\prime}={J}^{2}\rho_{{e}}-\frac{P\tilde{h}}{A}{\theta}\;,\;\theta(0)=\theta(L)=0,$ (6)
where $\kappa$ is the thermal conductivity, $\tilde{h}$ is the average convective heat transfer coefficient, and $P=b+2d$ is the perimeter of the beam’s cross-section. $J=I/A=V_{\mathrm{h}}/(\rho_{e}L)$ and $\rho_{e}$ represent the electric current density and the beam’s material resistivity, respectively. Assuming that $\kappa$ and $\rho_{e}$ are temperature-independent around the operating point, we can find the averaged solution of Eq. (6) as follows
$\overline{\theta}=\frac{V_{\text{h}}^{2}}{4\kappa\rho_{e}m^{2}}\left(\!1-\!\frac{\tanh(m)}{m}\right),\;\;m=\frac{{L}}{2}\sqrt{\frac{P\tilde{h}}{\kappa A}}.$ (7)
[…]
Table 3: Properties of single crystal silicon (correspond to $T=300$K).
Parameter Value
Young’s modulus, $E$ 169 GPa
[…]
Thermal conductivity, $\kappa$ $150$ W/(mK)
Vacuum permittivity, $\varepsilon_{0}$ $8.85\times 10^{-12}$ F/m
Resistivity, $\rho_{e}$ $0.02\;\Omega$ cm
[…]
Residual stress, $\sigma_{res}$$2.7$ MPa
In this study, we revisit the conclusions of the works kessler2018JMEMS ; krylov2018IEEE ; kessler2020flow ; kessler2021sampling and perform detailed experimental investigations of the influence of several additional, not addressed in the previous works, effects on the sensor. In the case of perpendicular (to the beam’s axis) flow, the drag force was found in krylov2018IEEE to influence the beam behavior. In the present work, the flow direction is varied, but most of the experimental results are obtained for the flow parallel to the beam axis ($\alpha=0^{\circ}$), where the contribution of the drag force is negligible. Therefore, we have not included the corresponding term into the analytical model.
2.5.2 Frequency response
Besides the quasi-static analysis, we also investigate the beam’s frequency response, which involves exploring the effects of beam oscillations after ST and SB transitions. This encompasses studying the fundamental frequency of the free vibrations, $f_{\text{res}}$, corresponding to the linearized version of Eq. (5):
$f_{\text{res}}=\frac{\omega_{0}}{2\pi}\sqrt{1-n+\frac{\left(3w_{m}^{2}+2\eta^{2}-6\eta w_{m}\right)}{16}-\frac{3\beta}{2\sqrt{(\gamma+\eta-w_{m})^{5}}}},$ (8)
where $\omega_{0}=4\pi^{2}\sqrt{EI_{yy}/(3\rho AL^{4})}$ is the fundamental frequency approximation of the initially straight unforced beam and $w_{m}(\beta,\overline{\theta})$ is the solution of the equilibrium equation Eq. (5). More information about this derivation can be found in krylov2010dynamic ; krakover2016displacement ; kessler2021sampling .
3 Results and discussion
In this section, we outline the results of measurements revealing various factors on sensor behavior during operation, such as the overheat ratio, flow loading, noise and vibrations. Finally, for the first time, we demonstrate the real-time beam shape during ST-SB transitions.
3.1 Joule heating effect
We present first the quantitative outcomes from various runs conducted at different Joule heating voltages ($V_{\mathrm{h}}$=0, 1, 1.33, and 1.66 V). In Figure 5(a, b), we illustrate the variations in $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ in response to Joule heating at different $V_{\mathrm{h}}$ values across a flow velocity range of 0 to 15 m/s. Notably, our results show that Joule heating induces a consistent offset in all curves. Interestingly, we find that sensitivity remains consistently high across all voltage levels, reflecting the robustness of the sensor’s performance. Note that for all these measurements, we used LDV with the inherent bias in $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ due to laser heating of the beam.
Figure 5: (a) Critical snap-through voltage ($V_{\mathrm{ST}}$) and (b) critical snap-back voltage ($V_{\mathrm{SB}}$) as a function of the flow velocity $U$ at different $V_{\mathrm{h}}$ levels. Dashed lines emphasize linear trend lines.
To provide a more comprehensive understanding of the impact of Joule heating voltage playing a key role in determining the flow sensor’s overheat ratio, we thoroughly investigate the beam’s response to a range of electrostatic voltages ($V_{\mathrm{ES}}$). We employ an analytical model, which is detailed in Sect. 2.5, to systematically compare the model’s predictions and the experimental results. This in-depth analysis allows us to assess how varying electrostatic voltages influence the sensor’s behavior, shedding light on its performance in different operational scenarios.
Figure 6(a) shows the analytical predictions of changes in the electrostatic voltage $V_{\mathrm{ES}}$ as a function of the normalized midpoint displacement $w_{m}$ of the beam.
Figure 6: (a) Modeled beam response to the electrostatic voltage $V_{\mathrm{ES}}$ at various heating voltages $V_{\mathrm{h}}$, with a few curves shown for clarity. Insets show the convex (before ST) and concave (after ST) configurations of the beam. (b) Comparison between experimental (symbols) and modeled (solid curve) voltage $\tilde{V}$ as a function of voltage $V_{\mathrm{h}}$. Error bars represent a standard error of the mean. (c) The critical voltage indicator ($\tilde{V}$) as a function of the flow velocity $U$ at different $V_{\mathrm{h}}$ levels. The model results are represented by lines (dotted, dash-dot, dashed, and solid for $V_{\mathrm{h}}$=0, 1, 1.33, and 1.66 V, respectively). Error bars represent a standard error.
The other limit point (local minimum, SB) on the curve corresponds to the SB voltage $V_{\text{SB}}$ at which the beam returns to its original state. It is important to note that bistability can manifest even without additional beam heating ($V_{\text{h}}=0$), attributed to the beam’s initial curvature.
As $V_{\mathrm{h}}$ increases, $V_{\text{ST}}$ increases, while the $V_{\text{SB}}$ decreases. Once a certain level of heating is reached, $V_{\text{SB}}$ reaches zero. The meaning is that the beam can stay in the switched configuration at zero actuating voltage. At even higher $V_{\mathrm{h}}$ , a force in the opposite (negative $z$) direction should be applied to switch the beam back. This behavior is known as the latching effect simitses2012dynamic .
To quantify the relationship between critical voltages $V_{\text{ST}}$ and $V_{\text{SB}}$, with respect to both flow velocity and Joule heating, we introduce a universal critical voltage indicator denoted as $\tilde{V}$, calculated as $\tilde{V}=(V_{\text{ST}}-V_{\text{SB}})/(V_{\text{ST}}+V_{\text{SB}})$.
This introduction of a universal indicator serves two key purposes. First, it consolidates the information into a single value, streamlining the analysis rather than dealing with separate ST and SB voltage values. Second, it attempts to disentangle the influence of the overheat ratio effect from the impact of varying ambient temperatures, a factor that can fluctuate in different experimental conditions.
Figure 6(b) compares the analytical model with the experimental results, illustrating the changes in $\tilde{V}$ as a function of $V_{\mathrm{h}}$. The model predictions are consistent with the experimental observations. Furthermore, it predicts the value of the heating voltage value ($V_{\text{h}}=2.2$ V) corresponding to latching, which was not reached in the experiment.
The Joule heating effect in terms of $\tilde{V}$ as a function of the flow velocity $U$ is shown in Figure 6(c). The figure also shows, represented by the different types of lines, the theoretical results obtained using Eq. (5). There is no perfect agreement between the model and the experiment, especially for the low heating case ($V_{\mathrm{h}}$=0 and 1 V). Note that the reduced-order model could not take into account the effect of velocity when the Joule heating is “off” ($V_{\mathrm{h}}$=0 V), but it is possible that there is a small temperature difference between the airflow and the microbeam. Nevertheless, the reduced-order model accurately captures the trends in the critical voltage change, indicating that the primary factor in velocity measurement is the change in average beam temperature, as suggested by the model. The graphs also show that all curves have nearly the same slope angle of 0.015 V/(m/s). This suggests that the beam overheat ratio is not a critical parameter for the sensor, which raises the possibility of significantly reducing energy consumption at the required sensitivity level.
3.2 Flow angle sensitivity
This section delves into the influence of flow angle on the sensor’s response. The sensor’s angle sensitivity is evaluated in the plane parallel to the sensor, as depicted in Figure 2(b, c), with the sensor’s long side aligned with the flow direction. The results of this impact, illustrated in Figure 7, reveal that the $\tilde{V}$ values remain consistent at low velocities. However, as flow velocities increase, deviations appear for different flow directions defined by the angles $\alpha=0^{\circ},15^{\circ}$, and $45^{\circ}$.
Figure 7: Influence of flow direction by changing the angle $\alpha^{\circ}$ ($V_{\mathrm{h}}$=1.33 V). Dashed lines represent visual trends. Error bars represent a standard error of the mean.
The effect of flow direction on the $\tilde{V}$ value is noticeable. Our sensor behaves similarly to miniature hot-wire sensors used for measuring turbulent flows, as shown in reference borisenkov2015multiarray . Due to its simpler manufacturing process and reliable angular response, the bistable microbeam has potential as a component in future multi-array flow sensors, capable of measuring all three flow velocity components.
3.3 Frequency response and sensitivity to vibrations
In Figure 4, it is evident that the sensors respond to flow velocity even when no LDV or Joule heating is involved. This suggests the existence of an additional mechanism associated with airflow in our experiment. Since the flow runs parallel to the flow sensor, it is unlikely that this effect is attributed to mean flow drag or lift forces, as these forces are negligible in this particular configuration.
We considered that the device might be sensitive to turbulent velocity fluctuations or chip vibrations. To investigate this, we initially determined the natural frequency of the beam’s vibration around various equilibrium configurations using LDV, following a similar approach to our previous study kessler2021sampling .
To achieve this, we introduced time-harmonic (AC) components to the steady (DC) electrostatic actuating voltage, resulting in the equation $V_{\text{ES}}(t)=V_{\text{DC}}+V_{\text{AC}}\sin{(2\pi f_{a}t)}$. Here, $V_{\text{DC}}$ represents the DC bias voltage, and $V_{\text{AC}}$ is the amplitude of the time-harmonic AC component, set at 0.5 V.
During the frequency up-sweep, the frequency $f_{a}$ increases linearly with time, ranging from 30 to 120 kHz. We determined the natural frequency of the beam’s vibration, denoted as $f_{n}$, for different equilibrium deflections determined by the DC bias voltage. The measured resonant frequency of the beam, when there is zero electrostatic voltage applied, is approximately 111 kHz.
In Figure 8, we present the normalized resonant frequency of the beam as a function of the normalized bias voltage.
Figure 8: Frequency of free vibrations around equilibrium as a function of the increasing/decreasing DC bias voltage. The baseline value $f_{0}$ represents the structure’s frequency in the initial unloaded configuration. The bias voltage is normalized by the snap-through voltage $V_{\mathrm{ST}}$. Experimental results for increasing and decreasing voltages are represented by circle and square markers, respectively. The solid and dash-dot lines represent the calculated increasing/decreasing curves.
The experimental data is depicted as circles and squares, while the calculated results, obtained using Eq. (8), are represented by solid and dash-dot lines. These dependencies are displayed for increasing and decreasing DC voltage ($V_{\text{DC}}$) values, covering the range encompassing the values at which ST and SB events (represented by $V_{\text{ST}}$ and $V_{\text{SB}}$) occur. Notably, the experimental findings align well with the results generated by the reduced-order model.
As expected, the vibration frequency significantly decreases near the ST and SB points. However, it is crucial to note that the experimental data does not reach absolute zero values, which would entail zero frequency and stiffness. This contrasts with the model’s predictions krylov2010dynamic .
The measured frequencies at the ST and SB branches reveal an exceptionally low beam stiffness in close proximity to the critical $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ points. Consequently, even a minor vibration and noise level can induce an ST transition at voltages lower than initially predicted.
We can derive consistent conclusions by examining the beam’s frequency response in the context of electrostatic actuation with a linear voltage ramp. In this analysis, we utilized the same voltage ramp signal, denoted as $V_{\mathrm{ES}}$, as previously detailed in Section 2.3, with a 60-second duration, allowing for a thorough examination of the beam’s vibrations. The LDV signal was sampled at a rate of 1 MHz.
Figure 9 illustrates the beam’s frequency response, particularly during the occurrences of snap-through (ST) and snap-back (SB) transitions. This power spectral density (PSD) spectrogram was generated through the application of the overlapped segment averaging spectral estimation (using the ’spectrogram’ Matlab function) with a segment length of $2^{18}$ samples and a 75% overlap.
The spectrogram reveals that the fundamental frequency closely mirrors the curves in Figure 8. Notably, a sharp change in the fundamental frequency is evident during the transition between equilibrium states, specifically during snap-through and snap-back. Moreover, we observe the emergence of a higher (second) harmonic vibration mode, surpassing the intensity of the first mode by a factor of 2.75. These results affirm our findings regarding the sensor’s high sensitivity in proximity to the beam’s equilibrium state. They underscore the fact that even a minor level of vibration or noise can trigger snap-through or snap-back transitions. These factors are known to hinder the experimental voltage-frequency dependency curves from reaching the theoretical zero level, as documented in references tilmans1994electrostatically ; hajjaj2017scalable ; torteman2019micro .
Figure 9: Frequency response of the beam vibrations in the form of a spectrogram during the voltage actuation ramp $V_{\mathrm{ES}}$ (60 s). The level of the PSD is normalized to the maximum value. The overheat ratio is $V_{\mathrm{h}}=$0 V. The line that corresponds to $1.8f_{0}$ is a regular noise signal that comes from the voltage amplifier.
To gain a deeper insight into this effect, we performed an experiment designed to emulate the influence of the noise on the ST and the SB voltages. Namely, we add an electrical white noise signal generated by the wave generator to the baseline DC electrostatic voltage $V_{\mathrm{ES}}$ used in all other experiments.
The noise amplitude, ranging from 0 to 5 V, should be viewed in the context of the base DC voltage, which extends from 0 to 200 V in the electrostatic voltage ramp signal. The measurements of $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ are carried out under conditions without flow ($U=0$ m/s) and are depicted in Figure 10.
Figure 10: Effect of noise on sensor performance ($V_{\mathrm{h}}$=0 V). The error bars represent a standard error of the mean.
As the amplitude of noise increases, its influence becomes apparent in the absolute value of $\tilde{V}$. This provides direct evidence that the beam exhibits remarkable sensitivity to fluctuations, including those induced by the flow, near the critical equilibrium point. Consequently, it is well-suited for sensing turbulent flows. This capability has the potential for applications beyond the measurement of average flow velocity. It can also be used to capture turbulent fluctuations, such as in shear stress measurements, particularly when the sensors are flush with walls. Notably, similar sensors based on hot-wire technology have been developed chew1998dynamic ; sturzebecher2001surface ; ghouila2019unsteady . Our findings open up new possibilities for applying bifurcation-based sensors in this field.
3.4 Dynamics of the ST transition
In contrast to previous studies kessler2020flow ; kessler2021sampling , where a single-point LDV system was employed to measure only the midpoint deflection, our approach in this study takes a significant step forward. We utilize the Micro System Analyzer (Polytec, MSA 600).
The MSA operates in a scanning mode, allowing us to capture the complete shape of the entire beam during the transient ST and SB events. To achieve this, the MSA records a 36 milliseconds long velocity signal from 56 points distributed along the length of the beam. We actuate the beam buckling with a linear ramp of $V_{\mathrm{ES}}$ ranging from 0 to 200 V at a frequency of 400 Hz. The displacement of the beam at each point is determined by integrating the velocity signal over time. Phase averaging is applied to calculate the average displacement at each measurement point on the beam, effectively eliminating free vibrations following ST and SB.
In Figure 11, we present snapshots of the beam’s shape $z(x)$ along the beam’s length $x$ during the snap-through and snap-back transitions.
Figure 11: Snapshots of beam deformation during snap-throught transition (a), and snap-back transition (b).
The arrows indicate the direction of the transitions, and the color coding reflects the progression from blue to magenta over time. We have marked positions at various arbitrary times, evenly distributed for clarity.
This represents a real-time beam profile measurement during the dynamic snap-through and snap-back transitions. Notably, we observe a slight asymmetry in the shape of $z(x,t)$ during both transitions concerning the midpoint, $z(x=0.5L)$. This implies that the transitions occur at lower $V_{\mathrm{ST}}$ and higher $V_{\mathrm{SB}}$ values than what would be anticipated from a symmetric first-mode transition. The presentation of these asymmetric snapshots of beam deformation during ST and SB events is a novel contribution, supporting the hypothesis presented in medina2012symmetry regarding the nature of this transition. Additionally, it provides insight into the shape of free vibrations following snap-through, in line with findings from song2021free .
Notably, during the beam transition, it is evident that the center of the beam lags behind the rest of the beam, with the exception of the edges, which remain static. Particularly, the sections between the center and the edges, around $0.40.2<x/L<$ 0.4 and $0.80.6<x/L<$ 0.8, lead the motion during a significant portion of the deformation. This shape likely reflects a combination of various modes and underscores the contribution of the second symmetric mode. This discovery encourages further exploration in this direction.
4 Summary and conclusions
This study delved into critical aspects of the sensor’s behavior. Firstly, we initiated by quantifying the “loading effect” induced by the laser Doppler vibrometer, shedding light on the impact of laser heating on voltage measurements.
Secondly, we investigated the beam’s response to varying levels of Joule heating. While the heating voltage $V_{\mathrm{h}}$ influenced the absolute values of $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ voltages, the device’s sensitivity remained notably consistent. Our analytical model supported these findings and revealed a latching effect with increasing Joule heating. These results suggest that the beam overheat ratio may not be as pivotal as assumed for most thermal MEMS sensors, promising substantial energy consumption reduction.
Thirdly, we investigated the influence of the flow angle on the performance of the beam. We concluded that the device performs similarly to previous hot-wire-based microsensors and can measure flow direction.
We conducted a detailed analysis of the beam’s response to fluctuations at different frequencies, typically present in various applications. Our results confirmed that this device, operating near critical transition points, is also sensitive to turbulent flow velocity fluctuations.
Finally, we obtained the first direct beam shape measurements during transitions and confirmed the theoretical predictions of asymmetric, higher-than-first-mode transitions. We can confirm that the single degree-of-freedom approximation of the single of beam needs further improvement.
Overall, this study lays the groundwork for designing a series of flow sensors and multi-sensor arrays for velocity and velocity gradient measurements using initially curved bistable microbeams. However, unanswered questions still require further investigation, such as the sensor’s response to out-of-plane flow directions, its performance in realistic environments, and the noise levels at which it can accurately measure low and high flow velocities for different applications.
Acknowledgments
The authors acknowledge the financial support of the Israeli Innovation Authority, NOFAR program and of RAFAEL LTD, Advanced Defense Systems. S. Krylov acknowledges the support of the Henry and Dinah Krongold Chair of Microelectronics. The authors would also like to thank Gal Spaer Milo, Grigori Gulitski, and Stella Lulinsky for their assistance.